Divide $128 into two parts such that one-third of the first part is less than
1
5
of the second part by
$8.

Answer by ewatrrr(24785)


Hi
$128 into two parts
Let x and ($128-x) represent the first part and second part respectively
If I am understanding Your question properly****
(1/3)x = (1/5)(128 -x) -$8 |multiplying each side by 15 so as all denominators= 1
5x = 3($128-x) - $8*15
5x = 3*128 - 3x - $8*15
8x = $264
x = $33, the first part. The second part = $95 ($128-$33)
CHECKING our Answer
(1/3)$33 = (1/5)$95 -$8
$11 = $19 - $8